OpenStudy (anonymous):

Change 4x – 3y = –15 to slope-intercept form. You must show all work. State the slope and y-intercept

OpenStudy (precal):

-3y=-15-4x y=(-15-4x)/-3 y=5+(4/3)x slopeis 4/3 y intercept is 5